Define electric displacement vector D vector and deduce Gauss's law for dielectrics


1
Expert's answer
2021-01-24T16:33:06-0500

Electric displacement vector D  is the charge per unit area that would be displaced across a layer of conductor placed across an electric field.  It is also known as flux density.

Now we know that according to gauss law,


"\Phi=\int E_0dA=\frac{q}{\epsilon_0}\\E_0=\frac{q}{A\epsilon_0}"

Where Φ is the flux , E0 is the electric field when no diectric is placed , A is the surface area, q is the charge and €o is permeability in vacuum.

Now let E be the electric filed when a dielectric is placed such that ,


"\epsilon=k\epsilon_0"

Where k is the dielectric constant and € is the permiability in the dielectric.

Now


"E=\frac{q}{A\epsilon}=\frac{q}{kA\epsilon_0}"